# Learning User Preferences Through Interaction for Long-Term Collaboration

### v1

@@ Abstract
Learning User Preferences Through Interaction for Long-Term Collaboration
@@ 1 Introduction
Throughout repeated interactions, humans naturally develop interpersonal relationships and adapt their behaviors to each other across various dimensions, including matching communication styles, establishing common ground, and refining mutual understanding (burgoon1995interpersonal; niederhoffer2002linguistic; wilkes1992coordinating; clark1991grounding; clark1986referring). This type of adaptation is particularly critical in human-AI collaboration, where users have interaction preferences (e.g., having key takeaways at the end of responses or receiving multiple options and their tradeoffs before making a decision) that influence how effectively they can understand and engage with agent responses (sweller1988cognitive; shi2025models). Rather than placing the cognitive burden on the user to repeatedly specify their preferences, conversational agents must be able to continuously learn and adhere to them over time (bickmore2005establishing; shi2025models).
@@
Recent works have introduced memory mechanisms to overcome the finite context window of Large Language Models (LLMs) and enable agents to dynamically store and retrieve user-specific information (chhikara2025mem0; wang2025mirix; yan2025memory). Evaluations of these systems generally focus on how accurately agents can recall information and answer questions about past interactions (xu-etal-2022-beyond; maharana2024evaluating; wu2025longmemeval). While such recall abilities are a necessary prerequisite, they do not capture all abilities needed for the intended purpose of these systems, which is to improve long-term interaction with users. In particular, it remains unclear whether agents can recognize what information is valuable for future interactions and leverage it effectively.
@@
To address this gap, we introduce MultiSessionCollab, a benchmark designed to evaluate how well conversational agents can learn user preference information and leverage it to meaningfully improve interactions during multi-session collaboration. The benchmark features user simulators that collaborate with the agent on problem-solving tasks across multiple sessions. Each user has distinct personas and interaction preferences that are derived from psychology, cognitive science, and human-computer interaction research, reflecting realistic human behaviors in collaborative settings (see taxonomy in Appendix A) (wood2007new; sweller2006worked; schwartz2002maximizing). The benchmark is designed so that users progress most effectively when agent responses align with their preferences, incentivizing agents to continuously learn and adhere to preferences. An overview of the benchmark is presented in Figure 1.
@@
Unlike prior memory systems that optimize for information recall, we develop long-term collaborative agents equipped with memory that learns user preferences to improve collaboration quality over time. As shown in Figure 1, agents reflect on interactions to identify user preference information that will be valuable for future interactions and update their memory accordingly. During subsequent sessions, the full memory is provided to the agent, and relevant parts are dynamically retrieved throughout the interaction to guide agent behavior towards being more preference-aligned.
@@
We also present a reinforcement learning (RL) framework that derives learning signals from user simulator behavior in MultiSessionCollab to train agents to update memory more effectively. We use Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) with an LLM-judge that rewards reflections for comprehensively capturing user preferences revealed during interaction.
@@
Extensive experiments across several models and five problem-solving tasks demonstrate that memory enables agents to continuously learn user interaction preferences and improve collaboration quality, yielding higher task success rates, more efficient interactions, and reduced user effort. Analyses of performance across sessions reveal continued improvement throughout the sessions, with the steepest gains occurring in early sessions before gradually stabilizing. Notably, agents that learn preferences through interaction are competitive with those given direct access to ground-truth user preferences, showing how memory captures richer information about preferences.
@@ 2 Related Work
Early works that evaluate LLMs in multi-session interactions focus on how well models can generate responses that are consistent with past interactions (xu-etal-2022-beyond; xu-etal-2022-long). More recently, benchmarks have shifted towards question-answering style evaluations that assess how well LLMs can remember information from past interactions (packer2023memgpt; maharana2024evaluating; wu2025longmemeval; hu2025evaluating; jiang2025know; zhao2025do). In contrast, our work evaluates the downstream impact of memory, incorporating abilities such as identifying what information will be useful in future interactions and leveraging it to improve collaboration.
@@
Providing LLMs with large contexts is computationally inefficient and can degrade performance, since LLMs struggle to effectively handle large amounts of information (shi2023large; liu2024lost). While one line of works try to improve LLM abilities to handle larger contexts (liu2025comprehensive), another line of works introduces memory to enable agents to store information from past experiences, and retrieve it when useful in future interactions (shinn2023reflexion; packer2023memgpt; zhong2024memorybank; suzgun2025dynamic; wang2025agent; ho2025arcmemo; chhikara2025mem0). More recent works introduce more sophisticated memory architectures, including those that leverage multi-agent systems (wang2025mirix) or temporal-aware knowledge graphs (rasmussen2025zep). While these works demonstrate the benefit of memory for question-answering tasks, we demonstrate that memory can improve downstream performance in multi-session collaboration settings.
@@
RL has been used to train agents to manage and utilize memory more effectively. Existing approaches use rewards based on question-answering correctness (zhou2025mem1; yan2025memory; wang2025mem; yu2025memagent). In contrast, we present an RL framework that derives rewards from user behavior during interaction. Specifically, we train agents to recognize what user preference information revealed during an interaction will be valuable for future sessions and update memory more effectively.
@@
Recent work has demonstrated that asking clarifying questions can enhance multi-turn interactions by helping agents better understand tasks, user intent, and preferences (zhang2025modeling; li2025alfa; li2025eliciting; andukuri2024stargate; wan2025enhancing; li2025personalized), with several works showing that this behavior is helpful for human-AI collaboration (wu2025collabllm; zhou2025sweet; wang2025beyond). These approaches focus on single-session interactions, which is appropriate for cold-start scenarios. However, as users increasingly engage with agents over multiple sessions, repeatedly asking the similar questions and eliciting preferences can become tedious and place unnecessary cognitive burden on users. Our work addresses this limitation by enabling agents remember user preference information across multiple sessions.
@@ 3.1 Collaborative Problem-Solving Session
In each session, a user simulator seeks to solve a problem with assistance from a conversational agent. Following wu2025collabllm; zhou2025sweet, only the user has access to the problem statement, and they start the conversation by providing an initial description of their problem. Over multiple turns, the agent asks clarifying questions to better understand the user’s task and provides explanations to help the user develop a solution. Conversations last until the user is satisfied and decides to terminate, or when the maximum conversation length of 10 user-agent turns has been reached. This design mirrors realistic collaborative scenarios, such as tutoring or debugging, and results in natural multi-turn interactions.
In such collaborative settings, agents must not only provide correct information, but also communicate it in ways that users can understand, apply, and learn from (shi2025models). We capture this by having users maintain a draft answer for their solution to the problem. The draft answer is never visible to the agent. It is initially empty, and is progressively updated as users receive assistance.
Crucially, users update their draft answer only when the agent provides helpful information in a manner that aligns with their interaction preferences. Information delivered in a way that violates the user’s preferences (e.g. using overly technical language when the user prefers simple explanations) hinders their ability to effectively understand and apply the information (shi2025models; sweller1988cognitive). In such cases, users explicitly communicate their preferences to the agent rather than updating their draft answer. We note that real-world users may be able to extract useful information even when responses do not fully align with their preferences. Our design choice of not updating the draft answer in such cases helps isolate preference adherence, enabling more direct evaluation of how well agents can learn and adapt to preferences across sessions.
@@ 3.2 User Simulator Design
First, we instantiate each user profile with a randomly selected persona from Persona Hub, a large-scale persona collection (ge2024scaling). Each persona exhibits unique knowledge, experiences, interests, personalities and professions, which translates to varying perspectives and diverse behaviors.
Each user profile is also assigned a random set of three interaction preferences that describe how they expect the agent to behave during collaboration. The preferences may specify specific communication styles (miehle2020estimating), learning approaches (chi1989self), or proactivity levels (horvitz1999principles). For instance, a user may prefer concise responses with key takeaways highlighted at the end, or expect detailed step-by-step explanations. Each preference is grounded in studies from psychology, cognitive science, and human-computer interaction, and reflects realistic human behavior in problem-solving settings. A complete taxonomy of the preferences and their sources are provided in Appendix A.
Finally, the effectiveness of our benchmark as an evaluation framework depends on user simulators that can reliably adhere to their user profiles and consistently enforce their preferences (mehri2025goal). To ensure this, we employ a structured reasoning process that explicitly considers whether (a) each preference has been satisfied and if any needs to be enforced, (b) the draft answer should be updated, and (c) the conversation should be terminated. This structure yields interpretable user behavior signals: we track which utterances enforce preferences, using this both to quantify user effort (Section 3.3) and to derive learning signals for our RL framework (Section 4.2). The system prompt is provided in Appendix B.
@@ 3.3 Evaluation Metrics
  • Task Success: the accuracy of the user’s final draft answer, which represents what the user learned from the interaction. Task success is a standard metric for collaboration (wu2025collabllm; wang2025beyond).
@@
These metrics represent collaboration quality: higher task success, lower user effort, and shorter conversations indicate more effective collaboration.
@@ 5 Experimental Setup
To rigorously evaluate conversational agents with MultiSessionCollab, we instantiate the benchmark with 100 user profiles and five problem-solving benchmarks that cover diverse domains: MATH-500 (hendrycks2021measuringmath), MATH-Hard (hendrycks2021measuringmath), LogiQA (liu2007logiqa), MMLU (hendrycks2021measuring), and MedQA (jin2021disease). We use Llama-3.3-70B-Instruct (dubey2024llama) for our user simulator and LLM-judge. Each user collaborates with the agent to solve 20 randomly sampled problems, with one problem per session and a maximum of 10 conversational turns per session. Across all benchmarks, this totals 10,000 collaborative sessions per agent, ensuring robust evaluation of agent performance during multi-session collaboration.
@@
To validate the reliability of MultiSessionCollab and understand the contribution of different components, we conduct a series of ablation studies that isolate the impact of various dimensions. We observe a drop in performance when moving from direct problem-solving to multi-turn interaction, which aligns with prior work showing that LLMs struggle to ask clarifying questions or maintain coherence across turns (wu2025collabllm; laban2025llms; zhou2025sweet; mehri2025goal). Performance decreases further when introducing user preferences, since users correct agent behavior during preference violations rather than focus on the task. Our complete analysis is provided in Appendix C. We also present a user study that confirms our findings generalize to real users in Section 7.
@@
We train Llama-3.1-8B-Instruct (dubey2024llama) and Qwen-2.5-7B-Instruct (qwen2025qwen25technicalreport) to generate session-level reflections given a conversation. The models are first initialized with supervised fine-tuning (SFT) as a cold start, which helps stabilizes RL optimization and leads to more effective performance (guo2025deepseek). Then, we conduct GRPO as described in Section 4.2. Details on hyperparameters and issues such as catastrophic forgetting are provided in Appendix D.
We evaluate several LLMs in MultiSessionCollab, including Llama-3.1-8B-Instruct, Qwen-2.5-7B-Instruct, gpt-oss-20b, and Llama-3.3-70B-Instruct (qwen2025qwen25technicalreport; openai2025gptoss120bgptoss20bmodel; dubey2024llama). Models are evaluated as standard conversational agents without memory, and also when equipped with memory. For Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models, we additionally report results after applying GRPO training to improve session-level reflections.
@@ 6 Results
Table 1 presents our experimental results on the MultiSessionCollab benchmark. Equipping conversational agents with memory enables them to learn preferences as they accumulate experience and proactively adhere to them without requiring users to repeatedly enforce them, reducing friction and allowing users to focus on the task. This generally results in improvements in collaboration quality: interactions become more efficient, require less user effort, and yield higher task success rates. We observe even further improvements when using our RL framework to train agents to capture more meaningful user preference information in session-level reflections and update memory effectively.
Analysis across different models.
Different models exhibit varying baseline performance, as shown in the first row for each model that reports the results for a standard conversational agent without memory. For instance, the Qwen family tends to achieve the highest performance on mathematical reasoning tasks. Regardless of the baseline performance, memory helps improve collaboration quality. How effectively a model can leverage memory depends on their abilities to capture information that will be valuable for future interactions through session-level reflections and leverage them during actual interactions to improve user experience.
RL for Improving Session-Level Reflections.
We apply our RL framework to train Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct to generate session-level reflections that effectively capture the preferences revealed during interactions. After training, both models are able to leverage the memory more effectively, with notable improvements across all metrics. In particular, while Qwen-2.5-7B-Instruct initially exhibited a $1.03\%$ decrease in task success when equipped with memory, after training, the same model achieves a $3.43\%$ improvement in task success. This demonstrates that the memory architecture’s effectiveness depends largely on reflection quality. Additionally, these results suggest that the way our RL framework uses learning signals derived from user simulator behavior offers a promising direction for improving agent memory management abilities.

### v2

# MultiSessionCollab: Learning User Preferences with Memory to Improve Long-Term Collaboration

@@ Abstract
MultiSessionCollab: Learning User Preferences with Memory to Improve Long-Term Collaboration
@@ 1 Introduction
Throughout repeated interactions, humans naturally develop interpersonal relationships and adapt their behaviors to each other across various dimensions, including matching communication styles, establishing common ground, and refining mutual understanding (Burgoon et al., 1995; Niederhoffer and Pennebaker, 2002; Wilkes-Gibbs and Clark, 1992; Clark and Brennan, 1991; Clark and Wilkes-Gibbs, 1986). This type of adaptation is particularly critical in human-AI collaboration, where users have interaction preferences (e.g., having key takeaways at the end of responses or receiving multiple options and their tradeoffs before making a decision) that influence how effectively they can understand and engage with agent responses (Sweller, 1988; Shi et al., 2025). Rather than placing the cognitive burden on the user to repeatedly specify their preferences, conversational agents must be able to continuously learn and adhere to them over time (Bickmore and Picard, 2005; Shi et al., 2025).
@@
Recent works have introduced memory mechanisms to overcome the finite context window of Large Language Models (LLMs) and enable agents to dynamically store and retrieve user-specific information (Chhikara et al., 2025; Wang and Chen, 2025; Yan et al., 2025). Evaluations of these systems generally focus on how accurately agents can recall information and answer questions about past interactions (Xu et al., 2022a; Maharana et al., 2024; Wu et al., 2025a). While such recall abilities are a necessary prerequisite, they do not capture all abilities needed for the intended purpose of these systems, which is to improve long-term interaction with users. In particular, it remains unclear whether agents can recognize what information is valuable for future interactions and leverage it effectively.
@@
To address this gap, we introduce MultiSessionCollab, a benchmark designed to evaluate how well conversational agents can learn user preference information and leverage it to meaningfully improve interactions during multi-session collaboration. The benchmark features user simulators that collaborate with the agent on problem-solving tasks across multiple sessions. Each user has distinct personas and interaction preferences that are derived from psychology, cognitive science, and human-computer interaction research, reflecting realistic human behaviors in collaborative settings (see taxonomy in Appendix A) (Wood and Neal, 2007; Sweller, 2006; Schwartz et al., 2002). The benchmark is designed so that users progress most effectively when agent responses align with their preferences, incentivizing agents to continuously learn and adhere to preferences. An overview of the benchmark is presented in Figure 1.
@@
Unlike prior memory systems that optimize for information recall, we develop more user-centric agents that are equipped with memory to learn user preferences and improve collaboration quality over time. As shown in Figure 1, agents reflect on interactions to identify user preference information that will be valuable for future interactions and update their memory accordingly. During subsequent sessions, the full memory is provided to the agent, and relevant parts are dynamically retrieved throughout the interaction to guide agent behavior towards being more preference-aligned.
@@
We also present a reinforcement learning (RL) framework that derives learning signals from user simulator behavior in MultiSessionCollab to train agents to update memory more effectively. We use Group Relative Policy Optimization (GRPO) (Shao et al., 2024) with an LLM-judge that rewards reflections for comprehensively capturing user preferences revealed during interaction.
@@
Extensive experiments across several models and five problem-solving tasks demonstrate that our memory enables agents to continuously learn user interaction preferences and improve collaboration quality, yielding higher task success rates, more efficient interactions, and reduced user effort. Analyses of performance across sessions reveal continued improvement throughout the sessions, with the steepest gains occurring in early sessions before gradually stabilizing. Notably, agents that learn preferences through interaction are competitive with those given direct access to ground-truth user preferences, showing how memory captures richer information about preferences.
@@ 2 Related Work
Early works that evaluate LLMs in multi-session interactions focus on how well models can generate responses that are consistent with past interactions (Xu et al., 2022a, b). More recently, benchmarks have shifted towards question-answering style evaluations that assess how well LLMs can remember information from past interactions (Packer et al., 2023; Maharana et al., 2024; Wu et al., 2025a; Hu et al., 2025; Jiang et al., 2025; Zhao et al., 2025). In contrast, our work evaluates memory for improving user interactions, incorporating abilities such as identifying useful information and actually leveraging it to improve collaboration.
@@
Providing LLMs with large contexts is computationally inefficient and can degrade performance, since LLMs struggle to effectively handle large amounts of information (Shi et al., 2023; Liu et al., 2024). While one line of works try to improve LLM abilities to handle larger contexts (Liu et al., 2025a), another line of works introduces memory to enable agents to store information from past experiences, and retrieve it when useful in future interactions (Shinn et al., 2023; Packer et al., 2023; Zhong et al., 2024; Suzgun et al., 2025; Wang et al., 2025c; Ho et al., 2025; Chhikara et al., 2025). More recent works introduce more sophisticated memory architectures, including those that leverage multi-agent systems (Wang and Chen, 2025) or temporal-aware knowledge graphs (Rasmussen et al., 2025). While these works demonstrate focus on memory for question-answering tasks, we present memory for improving user interactions during multi-session collaboration.
@@
RL has been used to train agents to manage and utilize memory more effectively. Existing approaches use rewards based on question-answering correctness (Zhou et al., 2025b; Yan et al., 2025; Wang et al., 2025b; Yu et al., 2025). In contrast, we derive rewards from user behavior in interactions. Specifically, we train agents to recognize what user preference information revealed during an interaction will be valuable for improving future sessions.
@@
Recent work has demonstrated that asking clarifying questions can enhance multi-turn interactions by helping agents better understand tasks, user intent, and preferences (Zhang et al., 2025; Li et al., 2025c, a; Andukuri et al., 2024; Wan et al., 2025; Li et al., 2025b), with several works showing that this behavior is helpful for human-AI collaboration (Wu et al., 2025b; Zhou et al., 2025a; Wang et al., 2025a). These approaches focus on single-session interactions, which is appropriate for cold-start scenarios. However, as users increasingly engage with agents over multiple sessions, repeatedly asking the similar questions and eliciting preferences can become tedious and place unnecessary cognitive burden on users. Our work addresses this limitation by enabling agents remember user preference information across multiple sessions.
@@ 3.1 Collaborative Problem-Solving Session
In each session, a user simulator seeks to solve a problem with assistance from a conversational agent. Following Wu et al. (2025b); Zhou et al. (2025a), only the user has access to the problem statement, and they start the conversation by providing an initial description of their problem. Over multiple turns, the agent asks clarifying questions to better understand the user’s task and provides explanations to help the user develop a solution. Conversations last until the user is satisfied and decides to terminate, or when the maximum conversation length of 10 user-agent turns has been reached. This design mirrors realistic collaborative scenarios, such as tutoring or debugging, and results in natural multi-turn interactions.
In such collaborative settings, agents must not only provide correct information, but also communicate it in ways that users can understand, apply, and learn from (Shi et al., 2025). We capture this by having users maintain a draft answer for their solution to the problem. The draft answer is never visible to the agent. It is initially empty, and is progressively updated as users receive assistance.
Crucially, users update their draft answer only when the agent provides helpful information in a manner that aligns with their interaction preferences. Information delivered in a way that violates the user’s preferences (e.g. using overly technical language when the user prefers simple explanations) hinders their ability to effectively understand and apply the information (Shi et al., 2025; Sweller, 1988). In such cases, users explicitly communicate their preferences to the agent rather than updating their draft answer. We note that real-world users may be able to extract useful information even when responses do not fully align with their preferences. Our design choice of not updating the draft answer in such cases helps isolate preference adherence, enabling more direct evaluation of how well agents can learn and adapt to preferences across sessions.
@@ 3.2 User Simulator Design
First, we instantiate each user profile with a randomly selected persona from Persona Hub, a large-scale persona collection (Ge et al., 2024). Each persona exhibits unique knowledge, experiences, interests, personalities and professions, which translates to varying perspectives and diverse behaviors.
Each user profile is also assigned a random set of three interaction preferences that describe how they expect the agent to behave during collaboration. The preferences may specify specific communication styles (Miehle et al., 2020), learning approaches (Chi et al., 1989), or proactivity levels (Horvitz, 1999). For instance, a user may prefer concise responses with key takeaways highlighted at the end, or expect detailed step-by-step explanations. Each preference is grounded in studies from psychology, cognitive science, and human-computer interaction, and reflects realistic human behavior in problem-solving settings. A complete taxonomy of the preferences and their sources are provided in Appendix A.
@@ 3.3 Evaluation Metrics
  • Task Success: the accuracy of the user’s final draft answer, which represents what the user learned from the interaction. Task success is a standard metric for collaboration (Wu et al., 2025b; Wang et al., 2025a).
@@
Higher task success, lower user effort, and shorter conversations indicate more effective collaboration.
@@ 5 Experimental Setup
To rigorously evaluate conversational agents with MultiSessionCollab, we instantiate the benchmark with 100 user profiles and five problem-solving benchmarks that cover diverse domains: MATH-500 (Hendrycks et al., 2021b), MATH-Hard (Hendrycks et al., 2021b), LogiQA (Liu et al., 2007), MMLU (Hendrycks et al., 2021a), and MedQA (Jin et al., 2021). We use Llama-3.3-70B-Instruct (Dubey et al., 2024) for our user simulator and LLM-judge. Each user collaborates with the agent to solve 20 randomly sampled problems, with one problem per session and a maximum of 10 conversational turns per session. Across all benchmarks, this totals 10,000 collaborative sessions per agent, ensuring robust evaluation of agent performance during multi-session collaboration.
@@
To validate the reliability of MultiSessionCollab and understand the contribution of different components, we conduct a series of ablation studies that isolate the impact of various dimensions. We observe a drop in performance when moving from direct problem-solving to multi-turn interaction, which aligns with prior work showing that LLMs struggle to ask clarifying questions or maintain coherence across turns (Wu et al., 2025b; Laban et al., 2025; Zhou et al., 2025a; Mehri et al., 2025). Performance decreases further when introducing user preferences, since users correct agent behavior during preference violations rather than focus on the task. Our complete analysis is provided in Appendix C. We also present a user study that confirms our findings generalize to real users in Section 7.
@@
We train Llama-3.1-8B-Instruct (Dubey et al., 2024) and Qwen-2.5-7B-Instruct (Qwen et al., 2025) to generate session-level reflections given a conversation. The models are first initialized with supervised fine-tuning (SFT) as a cold start, which helps stabilizes RL optimization and leads to more effective performance (Guo et al., 2025). Then, we conduct GRPO as described in Section 4.2. Details on hyperparameters and issues such as catastrophic forgetting are provided in Appendix D.
We evaluate several LLMs in MultiSessionCollab, including Llama-3.1-8B-Instruct, Qwen-2.5-7B-Instruct, gpt-oss-20b, and Llama-3.3-70B-Instruct (Qwen et al., 2025; OpenAI, 2025; Dubey et al., 2024). Models are evaluated as standard conversational agents without memory, and also when equipped with memory. For Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct models, we additionally report results after applying GRPO training to improve session-level reflections.
@@ 6 Results
Overall
TS (%) $\uparrow$ UE $\downarrow$ Len $\downarrow$
Llama-3.3-70B 41.78 2.98 15.96
  + Mem0 43.27↑1.49 2.19↓0.79 15.06↓0.90
  + memory (ours) 46.38↑4.60 1.96↓1.02 14.53↓1.43
Table 2: Comparison between our memory architecture and Mem0 on MultiSessionCollab. We report task success (TS), user effort (UE), and conversation length (Len) averaged across all tasks. Subscripts denote change relative to the baseline without memory.
Main Results.
Table 1 presents our experimental results on the MultiSessionCollab benchmark. Equipping conversational agents with memory enables them to learn preferences as they accumulate experience and proactively adhere to them without requiring users to repeatedly enforce them, reducing friction and allowing users to focus on the task. This generally results in improvements in collaboration quality: interactions become more efficient, require less user effort, and yield higher task success rates.
We observe even further improvements when using our RL framework to train to generate session-level reflections that effectively capture the preferences revealed during interactions for Qwen-2.5-7B-Instruct and Llama-3.1-8B-Instruct. After training, both models are able to leverage the memory more effectively, with notable improvements across all metrics. In particular, while Qwen-2.5-7B-Instruct initially exhibited a $1.03\%$ decrease in task success when equipped with memory, after training, the same model achieves a $3.43\%$ improvement in task success. This demonstrates that the memory architecture’s effectiveness depends largely on reflection quality. Additionally, these results suggest that the way our RL framework uses learning signals derived from user simulator behavior offers a promising direction for improving agent memory management abilities.
Comparison with Existing Memory.
Unlike existing memory architectures, which are primarily designed for information retrieval and question-answering, ours is designed to improve interactions with users. To demonstrate this, we compare against Mem0, a memory architecture that stores information in a vector database after each user-agent turn and dynamically retrieves them prior to response generation (Chhikara et al., 2025). We equip Llama-3.3-70B-Instruct with each memory architecture and present results averaged across all tasks in Table 2.
Our memory architecture outperforms Mem0 across all metrics. We attribute this to a fundamental difference in what each memory captures. Mem0’s memories tend to store factual information relevant to answering questions, whereas our reflections capture user-centric information about how to interact with the user. For example, a Mem0 memory is: "Evaluating the effectiveness of communication with the parents is a crucial step…" – a factual detail extracted from a previous response focused on answering the user’s question. In contrast, a memory from our architecture is: "The user prefers a narrative and engaging style, with multiple viable approaches to addressing the problem…". These results highlight that for agents to develop long-term relationships with users and improve interactions over time, memory architectures must be explicitly designed to be more user-centric.